On December 31, 2019, the Notes Receivable account at P. Davis Materials Corporation had a balance of $40,000, which represented a six-month, 5 percent note received from a customer on October 1.

During the week ended June 7, 2019, McCormick Media received $75,000 from customers for subscriptions to its magazine Modern Business. On December 31, 2019, an analysis of the Unearned Subscription Revenue account showed that half of the subscriptions were earned in 2019.

On November 1, 2019, Perez Realty Company rented a commercial building to a new tenant and received $36,000 in advance to cover the rent for six months. Upon receipt, the $36,000 was recorded in the Unearned Rent account.

On November 1, 2019, the Mighty Bucks Hockey Club sold season tickets for 50 home games, receiving $9,000,000. Upon receipt, the $9,000,000 was recorded in the Unearned Season Tickets Income account. At December 31, 2019, the Mighty Bucks Hockey Club had played 5 home games.


For each of the above independent situations, indicate the adjusting entry that must be made on the December 31, 2019, worksheet assuming no previous adjusting entries have been made during the year.
Answer #1

Adjusting Entries

Transaction

Accounts Title and Explanation

Debit

Credit

a

31-Dec

Interest Receivable

$          600.00

Interest Revenue

$          600.00

(Interest earned on note)

b

Unearned Subscription Revenue

$    37,500.00

Subscription Revenue

$    37,500.00

(Subscription earned)

c

Unearned rent

$    12,000.00

Rent Revenue

$    12,000.00

(Rent revenue earned for two months)

d

Unearned season ticket income

$ 900,000.00

Season ticket income

$ 900,000.00

Total

$ 950,100.00

$ 950,100.00

Working

Interest Receivable

=+((40000*5%)/12*6)/5*3

Interest Revenue

=+((40000*5%)/12*6)/5*3

(Interest earned on note)

Unearned Subscription Revenue

=75000/2

Subscription Revenue

=75000/2

(Subscription earned)

Unearned rent

=36000/6*2

Rent Revenue

=36000/6*2

(Rent revenue earned for two months)

Unearned season ticket income

=9000000/50*5

Season ticket income

=9000000/50*5